# The Evolution and Future Perspectives of Artificial Intelligence Generated Content

### v1

@@ VI-D2 GANs
As an unsupervised learning, the emergence of GANs is no less than an industrial revolution in the evolutionary history of AIGC. Meeting the increasing demand for generated content, GANs improve the quality and diversity of generated content while also expanding labeled datasets, which is crucial for training other models [4].
@@
Despite its potential, the original GAN framework faces challenges, such as training stability, limited controllability, high data requirements, and difficulties in achieving desired image quality. To address these limitations, researchers have proposed various extensions and refinements to the traditional GAN model. For instance, WGAN [104] and LSGAN [105] improve training stability by replacing the binary cross-entropy loss function with Wasserstein distance and least squares loss, respectively. These modifications address issues like mode collapse in WGAM and gradient vanishing in LSGAN. To enhance controllability in producing outputs with specific attributes, researchers have explored approaches such as manipulating latent space as seen in InfoGAN [106] and introducing conditional variables in CGAN [107]. To reduce GAN’s dependency on large labeled datasets, Pix2Pix [108] uses paired datasets in supervised learning while CycleGAN [109] and StarGAN [110] employ unsupervised learning to avoid the need for paired data. In a similar fashion, researchers have sought to improve output resolution and diversity by scaling up the model capacity in BigGAN [111] or introducing a control mechanism to manipulate image features for greater diversity in StyleGAN [112].
@@ VI-E Example
Figure 12: Postional embeddings.
@@
The deep learning phase marks the official culmination of AIGC technology. The vast potential offered by neural networks has inspired researchers to achieve remarkable advancements. However, the amount of training required by neural networks and its associated high cost both in time and resources, present a significant challenge for research to further progress. The work on fine-tuning and reusing pre-trained models through transfer learning is emerging.
@@ VII Transfer Learning and Pre-trained Models
It is worth noting that in addition to being categorized based on the type of knowledge transferred, TL can also be categorized from other perspectives. For instance, based on the nature of tasks, TL can be classified into isomorphic transfer[113], where source and target tasks are similar and heterogeneous transfer, where they are different. Similarly, TL can be divided into supervised [114], semi-supervised [115], and unsupervised [116] based on the level of supervision in the source and target domains. For clarity, this paper categorizes all TLs based on the type of transferred knowledge and provides details on their differences and commonalities in Sub-section VII.A.
@@ VII-A4 Model-Based TL
Model-based TL focuses on reusing the entire model architecture and its learned parameters from source tasks for targeted tasks [121]. Typically, the reusable model is pre-trained on a large, related dataset. To adapt it to a new application, the early layers are frozen to retain general features while the later layers are fine-tuned for task-specific adaption. The number of layers retrained depends on the complexity of the target task. For example, a small dataset of manufacturing defect images can be used to fine-tune a pre-trained ResNet, allowing it to recognize defect-specific patterns [122]. By fine-tuning or retraining partial layers of the model, the approach significantly improves the model’s efficiency and target tasks [123].
VII-B Example
The emergence of the Large Language Model (LLM) provides a solid foundation for advancing TL techniques. Their versatility allows them to be applied across the aforementioned TL categories, enabling broader and more effective applications across diverse domains. Among the currently popular LLMs, such as OpenAI’s GPT[124]][1], Google’s Gemini (Bard)[125], Meta’s Llama series[2], and Anthropic’s Claude[126], Meta’s Llama series stands out as a well-known example. The latest model, Llama 3, offers high capacity and open-source accessibility, making it easy for fine-tuning and retraining, which meets various application needs. Therefore, we choose Llama 3 to exemplify the TL process in responding to the same prompt: ”Generate a research question with the key words: artificial intelligence, healthcare, and ethical implications”.
@@ VII-B Example
To achieve ”Transfer”, Llama 3 has two important transfer steps: pre-train and fine-turning. Llama 3 is pre-trained on a massive dataset. The process allows it to grasp grammar, semantics, and even some world knowledge. This aligns with Model-Based TL, where the transformer architecture itself facilitates knowledge transfer to various downstream tasks. During pre-training, Llama 3 also extracts transferable feature representations, such as token embeddings and positional encodings (RoPE[127]). These features capture universal linguistic patterns, laying the groundwork for subsequent fine-tuning.
@@
While pre-training provides a foundation, the fine-tuning step, on the other hand, achieves a specialization in the problematic research area. Llama 3 usually hones in on datasets related to AI and ethics. The fine-tuning process refines its vast knowledge “library” by deepening its understanding of these areas, and also extends to the unfamiliar healthcare field. After selecting a proper framework like LoRA[128] or QLoRA[129], the source model is adapted to the new target field with our fed healthcare-related dataset. This process also grasps the nuances, such as the specific applications of AI in healthcare (e.g., diagnosis and treatment), and the ethical challenges they pose (e.g., bias and privacy issues). In addition, fine-tuning strengthens connections within the knowledge base, also known as relationship-based transfer, helping Llama 3 identify and prioritize relevant information. Compared to previous stages’ examples, Llama 3 learns to recognize the intersection of AI, healthcare, and ethics, allowing it to generate research questions that focus precisely on these critical and interdisciplinary topics.
@@
When prompted, Llama 3 generates a response using a combination of pre-trained and fine-tuned knowledge. First, it identifies key concepts in the request, such as “artificial intelligence,” “healthcare,” and “ethical implications,” and considers them as core elements. It then formulates a question based on its understanding of the structure of the research question and the relationship between these keywords. The specific steps in the generation process are very similar to deep learning’s decoder, but at the same time the optimization algorithms such as RoPE[127] and GQA[130] are used to ensure the efficiency of the operation of the huge language model. This process makes it possible to ask a relevant and insightful question such as “What are the ethical implications of using AI for personalized healthcare treatment?
@@
In summary, TL provides a framework for knowledge adaptation and application. It enriches the deep learning ecosystem by reusing existing systems. It provides a way to overcome the limitations of deep learning. By transferring and adapting pre-trained and learned models to cope with new requirements.
VIII AIGC limitations and challenges
@@
AIGC technology has been under development for quite some time, and it is now accessible and usable by a wide range of users, demonstrating undeniable potential. However, there are several limitations and challenges that AIGC presents, which we must overcome to advance its development and applications. We categorize them into two primary directions: data-related issues and model structure limitations, discussed in Subsections A and B, respectively.
@@ VIII-A Data-related limitations and challenges
As elaborated earlier, data is the key for AIGC regardless of which types of methods are used for content creation. The greater the volume and dimensions of data, the more knowledge methods can gather, and the better content to be created. Dependency on data is also the source of many limitations and challenges.
First, the quality of AIGC is constrained by how much data the methods have “seen” and understand. The availability and accessibility of data remain the barriers that current AIGC struggles to overcome while research advancements in areas, such as transfer learning outlined in Section VII, few-short, and zero-short learning [131], strive to reduce this reliance.
Inevitably, biases, incorrect or even inappropriate content[132] might be present in the available training data, causing AIGC systems to perpetuate misinformation or unfair results. Therefore, data augmentation, such as data filtering[133], resampling[134], adversarial training[135], synthetic data generation[4], and data validation[136], is essential for diverse and balanced training datasets. Additionally, establishing ethical guidelines and legal frameworks [137] and conducting regular bias audits, although out of the scope of this paper, are also crucial to mitigating risks and ensuring fairness in AIGC.
Training data only presents past information, which can cause generated content to appear right but is in fact wrong, leading to hallucinations. That is because AIGC models can’t access up-to-date knowledge [138] during training. Recent development in Retrieval Augmented Generation (RAG) presents a promising solution [139]. By combining the strengths of retrieval-based and generation-based approaches, models’ generative abilities are enhanced – extremely critical for applications, such as in fields like law and medicine, where real-time factual accuracy is essential.
@@ VIII-B Model-structure-related limitations and challenges
As AIGC evolves from early rule-based systems to deep learning, the size and sophistication of models have been increasing dramatically to improve content generation. The growth in model structures demands more processing power, memory, and storage to handle vast amounts of data necessary for both training and real-time operations. Such large capital investment in time and power limits the widespread adoption of AIGC. Although there are no immediate solutions, some potential long-term directions include advancements in hardware accelerators (e.g., GPUs and TPUs [140]), memory storage technologies, and renewable energy to lower computational resource consumption and costs. Recent efforts in model optimization[141] and decentralized training across multiple devices with local data [142] present promising avenues for reducing computational burdens.
Deep learning is often viewed as a “black box” [143]. This lack of transparency makes it difficult for users to understand and trust its processes and outcomes fully. As a result, deep-learning-based AIGC faces similar trustworthiness issues, and the reduced confidence in generated content limits its broader acceptance. However, ongoing research on explainable AI [144][145] and trustworthy AI[146] is offering promising solutions to address these challenges.
@@
Finally, ensuring coherence and relevance in AIGC while preserving its creativity is a significant challenge and a long-term goal for its development. While there are no definitive solutions yet, we believe that human-in-the-loop AIGC is the way forward. As we progress toward a human-centric industry 5.0, improved human-machine collaboration will lead to higher-quality and more innovative generated content, making it an ideal approach to achieve AIGVC’s essential objectives.
@@ VIII-C Beyond technique issues
In addition to these core technical limitations, there are ethical and policy-related concerns. AIGC systems can be exploited for malicious activities, such as scams and phishing[150], while issues like copyright disputes[151], data privacy[152], and academic integrity continue to emerge. The ability of AIGC to amplify biases in training data, compromise personal information, and blur the lines between originality and plagiarism poses significant challenges. These ethical and technical challenges highlight the need for strong legal frameworks, ethical guidelines, and technological safeguards to ensure responsible AI deployment and development aligned with societal values. Collaboration among industry, academia, and regulators is essential to establish principles of fairness, accountability, and transparency while remaining adaptable to technological advancements[137].

### v2

@@ VI-D1 VAEs
Building on the success of VAEs, recent applications such as Stable Diffusion [98] have emerged. These applications extensively utilize diffusion models, which generate high-quality images through an iterative denoising process. A generative diffusion model learns data distribution by gradually adding noise during training and reversing the process to denoise and generate samples. Early diffusion models, however, suffer from drawbacks such as high computational costs and slow generation speeds. By integrating the latent-space compression features of VAEs, Stable Diffusion overcomes these limitations, achieving higher-resolution, enhanced efficiency, greater diversity, and improved textual controllability in generated outputs.
@@ VI-D2 GANs
As an unsupervised learning, the emergence of GANs is no less than an industrial revolution in the evolutionary history of AIGC. Meeting the increasing demand for generated content, GANs improve the quality and diversity of generated content while also expanding labeled datasets, which is crucial for training other models.
@@
Despite its potential, the original GAN framework faces challenges, such as training stability, limited controllability, high data requirements, and difficulties in achieving desired image quality. To address these limitations, researchers have proposed various extensions and refinements to the traditional GAN model. For instance, WGAN [99] and LSGAN [100] improve training stability by replacing the binary cross-entropy loss function with Wasserstein distance and least squares loss, respectively. These modifications address issues like mode collapse in WGAN and gradient vanishing in LSGAN. To enhance controllability in producing outputs with specific attributes, researchers have explored approaches such as manipulating latent space as seen in InfoGAN [101] and introducing conditional variables in CGAN [102]. To reduce GAN’s dependency on large labeled datasets, Pix2Pix [103] uses paired datasets in supervised learning while CycleGAN [104] and StarGAN [105] employ unsupervised learning to avoid the need for paired data. In a similar fashion, researchers have sought to improve output resolution and diversity by scaling up the model capacity in BigGAN [106] or introducing a control mechanism to manipulate image features for greater diversity in StyleGAN [107].
@@ VI-E Example
Figure 14: Positional embeddings.
@@
The deep learning phase marks the official culmination of AIGC technology. The vast potential offered by neural networks has inspired researchers to achieve remarkable advancements. However, the amount of training required by neural networks and their associated high cost both in time and resources, present a significant challenge for research to further progress. The work on fine-tuning and reusing pre-trained models through transfer learning is emerging.
@@ VII Transfer Learning and Pre-trained Models
It is worth noting that in addition to being categorized based on the type of knowledge transferred, TL can also be categorized from other perspectives. For instance, based on the nature of tasks, TL can be classified into isomorphic transfer[108], where source and target tasks are similar, and heterogeneous transfer, where they are different. Similarly, TL can be divided into supervised [109], semi-supervised [110], and unsupervised [111] based on the level of supervision in the source and target domains. For clarity, this paper categorizes all TLs based on the type of transferred knowledge and provides details on their differences and commonalities in Subsection VII.A.
@@ VII-A4 Model-Based TL
Model-based TL focuses on reusing the entire model architecture and its learned parameters from source tasks for targeted tasks [116]. Typically, the reusable model is pre-trained on a large, related dataset. To adapt it to a new application, the early layers are frozen to retain general features while the later layers are fine-tuned for task-specific adaptation. The number of layers retrained depends on the complexity of the target task. For example, a small dataset of manufacturing defect images can be used to fine-tune a pre-trained ResNet, allowing it to recognize defect-specific patterns [117]. By fine-tuning or retraining partial layers of the model, the approach significantly improves the model’s efficiency and target tasks [118].
VII-B Modern Pre-trained Models: Large Language Models (LLMs)
With the transfer objective determined, the next key step in transfer learning is the choice of a suitable pre-trained model. These models, trained on large datasets to capture general patterns within a source domain, enable the transfer of this knowledge to a target domain. Pre-trained models are utilized across various domains, including computer vision, speech processing, and multimodal applications. This subsection focuses specifically on Large Language Models (LLMs), given their growing importance and popularity. Built upon the Transformer architecture, LLMs feature structural variations tailored to different purposes, affecting how they process information and what tasks they excel at.
VII-B1 Decoder-Only Models
Decoder-based models process information in a strictly unidirectional (left-to-right) manner using an auto-regressive attention mechanism. This configuration enables such a model to predict each token based solely on previously generated tokens, making it highly effective for generative tasks such as open-ended text generation, dialogue, and code completion. Models, such as GPT-4[1], Claude [119], PaLM[120], and LLaMA [2], follow this design and are well-suited for scenarios requiring fluent and coherent output.
VII-B2 Encoder-Only Models
As illustrated in Section VII.C, the encoder in the Transformer architecture is responsible for understanding and extracting features from the input using bidirectional self-attention. By considering both left and right context simultaneously when encoding each token, this type of model develops a deep understanding of the input as a whole. Encoder-only models such as BERT[94] excel in comprehension tasks such as sentence classification, named entity recognition, extractive question answering, and retrieval.
VII-B3 Encoder–Decoder Models
As the name suggests, the Encoder-Decoder models combine both components of the Transformer architecture. Employing bidirectional self-attention in the encoder and unidirectional attention in the decoder, the models not only comprehensively capture the meaning of an input sequence but also ensure fluent and context-aware token generation. Models such as T5 [121] and BART [122] follow this structure design and are especially well-suited for tasks that require complex input-output mappings, such as machine translation, summarization, and text-to-text generation.
VII-C Example
The emergence of LLM provides a solid foundation for advancing TL techniques. In this subsection, we use Llama 3 to exemplify the TL process in responding to the same prompt: ”Generate a research question with the keywords: artificial intelligence, healthcare, and ethical implications.”.
@@ VII-C Example
To achieve ’Transfer’, Llama 3 has two important transfer steps: pre-training and fine-tuning. Llama 3 is pre-trained on a massive dataset. The process allows it to grasp grammar, semantics, and even some world knowledge. This aligns with Model-Based TL, where the transformer architecture itself facilitates knowledge transfer to various downstream tasks. During pre-training, Llama 3 also extracts transferable feature representations, such as token embeddings and positional encodings (RoPE[123]). These features capture universal linguistic patterns, laying the groundwork for subsequent fine-tuning.
@@
While pre-training provides a foundation, the fine-tuning step, on the other hand, achieves a specialization in the problematic research area. Llama 3 usually hones in on datasets related to AI and ethics. The fine-tuning process refines its vast knowledge “library” by deepening its understanding of these areas, and also extends to the unfamiliar healthcare field. After selecting a proper framework like LoRA or QLoRA[124], the source model is adapted to the new target field with our fed healthcare-related dataset. This process also grasps the nuances, such as the specific applications of AI in healthcare (e.g., diagnosis and treatment), and the ethical challenges they pose (e.g., bias and privacy issues). In addition, fine-tuning strengthens connections within the knowledge base, also known as relationship-based transfer, helping Llama 3 identify and prioritize relevant information. Compared to previous stages’ examples, Llama 3 learns to recognize the intersection of AI, healthcare, and ethics, allowing it to generate research questions that focus precisely on these critical and interdisciplinary topics.
@@
When prompted, Llama 3 generates a response using a combination of pre-trained and fine-tuned knowledge. First, it identifies key concepts in the request, such as “artificial intelligence,” “healthcare,” and “ethical implications,” and considers them as core elements. It then formulates a question based on its understanding of the structure of the research question and the relationship between these keywords. The specific steps in the generation process are very similar to deep learning’s decoder, but at the same time, the optimization algorithms such as RoPE [123] and GQA [125] are used to ensure the efficiency of the operation of the LLM. This process makes it possible to ask a relevant and insightful question, such as “What are the ethical implications of using AI for personalized healthcare treatment?”.
@@
In summary, TL provides a framework for knowledge adaptation and application. This enriches the deep learning ecosystem by reusing existing systems. It provides a way to overcome the limitations of deep learning by transferring and adapting pre-trained and learned models to cope with new requirements.
VIII AIGC limitations, challenges, and future trends
@@
AIGC technology has been under development for quite some time, and it is now accessible and usable by a wide range of users, demonstrating undeniable potential. However, AIGC presents several limitations and challenges that pose barriers to its continued advancement. In this section, we examine these issues and explore future prospects for overcoming them. Specifically, we categorize the challenges into two primary directions: data-related issues and model structure limitations, discussed in Subsections A and B, respectively.
@@ VIII-A Data-related limitations and challenges
As elaborated earlier, data is the key for AIGC regardless of which types of methods are used for content creation. The greater the volume and dimensions of data, the more knowledge methods can gather, and the better the content to be created. Dependency on data is also the source of many limitations and challenges.
First, the quality of AIGC is constrained by how much data the methods have “seen” and understand. The availability and accessibility of data remain the barriers that current AIGC struggles to overcome, while research advancements in areas, such as transfer learning outlined in Section VII, few-shot[126], and zero-shot learning [127], strive to reduce this reliance. These methods are specifically designed to help generative models sustain high performance with minimal new data. In this context, few-shot learning allows models to generalize from a small number of labeled examples, while zero-shot learning supports task execution without any task-specific training data. In AIGC, these techniques enable applications such as summarization, question answering, and image synthesis, especially in domains with scarce or no annotated data.
Inevitably, biases, incorrect or even inappropriate content [128] might be present in the available training data, causing AIGC systems to perpetuate misinformation or unfair results. Therefore, data augmentation, such as data filtering [129], resampling [130], adversarial training[131], synthetic data generation [4], and data validation [132], is essential for diverse and balanced training datasets. Additionally, emerging strategies for improving data quality and aligning LLMs with human preferences, such as Direct Preference Optimization (DPO) and OpenAI’s Reinforcement Learning from Human Feedback (RLHF), merit further investigation. Finally, establishing ethical guidelines and legal frameworks [133] and conducting regular bias audits, although out of the scope of this paper, are also crucial to mitigating risks and ensuring fairness in AIGC.
Training data only presents past information, which can cause generated content to appear right but is in fact wrong, leading to hallucinations. That is because AIGC models can’t access up-to-date knowledge [134] during training. Recent development in Retrieval Augmented Generation (RAG) presents a promising solution [135]. By combining the strengths of retrieval-based and generation-based approaches, models’ generative ability is enhanced, extremely critical for applications, such as in fields like law and medicine, where real-time factual accuracy is essential.
Overall, data is still the most fundamental bottleneck in the development of AIGC, driving its future progress along two complementary directions. On the one hand, there is a growing effort to develop more data-efficient models that can achieve strong generative performance with fewer resources. Approaches such as transfer learning, few-shot, and zero-shot learning represent key strategies in this direction. On the other hand, expanding and improving the quality of training data is equally critical. Beyond real-world datasets, synthetic data, as a renewable “fuel” for AI, is gaining increasing attention for training and fine-tuning generative models. However, this trend also reveals new challenges: Can such data truly sustain the AIGC ecosystem over the long term? Without proper safeguards, synthetic data risks amplifying existing biases [136], reinforcing hallucinations [137], and even leading to “Habsburg AI” effects, such as performance degradation and systemic collapse [138]. To address these challenges, future research should focus on better integration of synthetic and real-world data with stronger human oversight, both during data curation and model alignment. This includes enhanced tools for data auditing and traceability [139], as well as human-in-the-loop methods such as DPO that align model outputs with human intent. These efforts are ultimately essential to building high-quality, low-bias, and highly controllable datasets.
@@ VIII-B Model-structure-related limitations and challenges
As AIGC evolves from early rule-based systems to deep learning, the size and sophistication of models have been increasing dramatically to improve content generation. The growth in model structures demands more processing power, memory, and storage to handle vast amounts of data necessary for both training and real-time operations. Such a large capital investment in time and power limits the widespread adoption of AIGC. Although there are no immediate solutions, some potential long-term directions include advancements in hardware accelerators (e.g., GPUs and TPUs [140]), memory storage technologies, and renewable energy to lower computational resource consumption and costs. Recent efforts in model optimization[141] and decentralized training across multiple devices with local data [142] present promising avenues for reducing computational burdens. For example, the Mixture of Experts(MoE) architecture—a recent variation of LLMs—introduces a scalable and efficient mechanism to reduce computation by activating a few lightweight ”expert” subnetworks for each input via a dynamic gating mechanism. In a similar fashion, within the context of LLMs, parameter-efficient fine-tuning methods such as LoRA allow models to be adapted to new tasks by training only a small subset of parameters to cut down memory and compute requirements.
Deep learning is often viewed as a “black box” [143]. This lack of transparency makes it difficult for users to understand and trust its processes and outcomes fully. As a result, deep-learning-based AIGC faces similar trustworthiness issues, and the reduced confidence in generated content limits its broader acceptance. However, ongoing research on explainable AI[144] and trustworthy AI[145] is offering promising solutions to address these challenges.
@@
Finally, ensuring coherence and relevance in AIGC while preserving its creativity is a significant challenge and a long-term goal for its development. While there are no definitive solutions yet, we believe that human-in-the-loop AIGC is the way forward. As we progress toward a human-centric industry 5.0, improved human-machine collaboration will lead to higher-quality and more innovative generated content, making it an ideal approach to achieve AIGC’s essential objectives.
The three main challenges of AIGC model structures—scalability, opacity, and vulnerability—continue to impact its broader adoption. Future research should shift focus from merely increasing model size and power to enhancing controllability, explainability, and robustness. Promising directions include modular architectures like MoE for efficient scaling, and parameter-adaptive methods, such as LoRA, for streamlined fine-tuning. Integrative approaches like neuro-symbolic integration [149], though still emerging in the AIGC domain, also hold potential to move generative systems beyond black-box behavior towards more explainable architectures by combining neural networks with interpretable symbolic reasoning. In other words, the creation of lightweight, energy-efficient, and general-purpose frameworks is crucial to make AIGC systems accessible to everyone.
@@ VIII-C Beyond technique issues
In addition to these core technical limitations, there are ethical and policy-related concerns. AIGC systems can be exploited for malicious activities, such as scams and phishing[150], while issues like copyright disputes[151], data privacy[152], and academic integrity continue to emerge. The ability of AIGC to amplify biases in training data, compromise personal information, and blur the lines between originality and plagiarism poses significant challenges. These ethical and technical challenges highlight the need for strong legal frameworks, ethical guidelines, and technological safeguards to ensure responsible AI deployment and development aligned with societal values. Collaboration among industry, academia, and regulators is essential to establish principles of fairness, accountability, and transparency while remaining adaptable to technological advancements.
In summary, AIGC raises a range of ethical and policy challenges, including misuse, privacy risks, copyright issues, biases, and academic integrity concerns. These problems highlight the urgent need for legal rules[133], ethical standards[153], and technical safeguards[154] that grow alongside the technology. To support responsible development, principles such as fairness, transparency[155], and accountability must be built into AIGC systems from the start. Addressing these challenges in a timely and proactive way will help ensure that AIGC benefits society while reducing risks. Strong governance, aligned with technical progress, is essential for building trust and ensuring that generative AI is used safely and fairly.